Two sand hauling trucks one that holds 12 tons and one that holds 15 tons make a total of 20 trips to haul 273 tons of sand. How many trips does each truck make?
Answer by checkley71(8403)

15X+12(20-X)=273
15X+240-12X=273
3X=273-240
3X=33
X=11 TRIPS FOR THE 15 TON TRUCK.
20-11=9 TRIPS FOR THE 12 TON TRUCK.
PROOF
15*11+12*9=273
165+108=273
273=273